# A Dynamical Theory of Overfitting and Generalization in Large Two-Layer Networks

### v2

@@ 2.4 Training on data with latent structure: mean field initialization
Figure 4: Training dynamics under a single-index model. We set $h(q)=\widehat{\varphi}(q)=(9/10)q+q^{3}/6$, $\tau=0.3$ and $\alpha=0.3$, under mean field initialization. Left: second-layer weights. Right: train and test error. Symbols are empirical results for SGD with actual two-layer neural networks with $d=200$, $n=\alpha md$. Lines correspond to predictions from the DMFT (on the right, continuous: train error; dashed: test error).
We initialize $a(0)=a_{0}$, independent of $m$ and let second layer weights evolve. Our DMFT analyisis predicts two dynamical regimes for large $m$. We will refer to them as ‘first’ and ‘third regime’ for consistency with other settings (see SM). For an illustration, we refer to Figs. 4 and 5.
@@
Namely, $\lim_{t\to\infty}a(t)={a}^{\mbox{\tiny\rm mf1}}(t)$, $\lim_{t\to\infty}{\boldsymbol{v}}(t)={{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}(t)$, and DMFT reduces to a system of $k+1$ ordinary differential equations for the $k+1$ scalar variables $({a}^{\mbox{\tiny\rm mf1}}(t),{{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}(t))$
@@
An explicit formula for ${e}^{\mbox{\tiny\rm mf1}}(t)$ is derived in the SM. In the case $k=1$ and $\widehat{\varphi}(z)=h(z)$, we have that ${a}^{\mbox{\tiny\rm mf1}}=1$, ${v}^{\mbox{\tiny\rm mf1}}=1$ is a fixed point of Eq. (2.9), and indeed the only fixed point with ${v}^{\mbox{\tiny\rm mf1}}>0$. If $h^{\prime}(0)>0$, then, we have $({a}^{\mbox{\tiny\rm mf1}}(t),{v}^{\mbox{\tiny\rm mf1}}(t))\to(1,1)$ as $t\to\infty$, and therefore test and train error converge to the Bayes error ${e}^{\mbox{\tiny\rm mf1}}(t)\to\tau^{2}/2$. This is of course significantly smaller than the test error achieved with lazy initialization. The separation between lazy and mean-field initialization is expected because feature learning takes place in the mean field regime.
Third dynamical regime: $t=\Omega(m)$. Computing the local stability of DMFT solutions around the mean field asymptotics (see SM) suggests that the latter breaks down for $t=\Theta(m)$. For $t\gtrsim m$, we observe that the second layer weights grow to achieve $a(t)\asymp\sqrt{m}$, the projection onto the latent space decreases to ${\boldsymbol{v}}(t)\asymp 1/\sqrt{m}$, and train and test error diverge, eventually achieving $e_{\mbox{{\rm\footnotesize tr}}}(t)\approx 0$ and test error significantly larger than the Bayes error achieved earlier. We refer to this phenomenon as ‘feature unlearning.’
@@
We observe that the limit curves $(\gamma,{{\varepsilon}}^{\mbox{\tiny\rm mf}}_{\mbox{{\rm\footnotesize tr}}}(\gamma))$, $(\gamma,{{\varepsilon}}^{\mbox{\tiny\rm mf}}_{\mbox{\rm\footnotesize ts}}(\gamma))$, match closely asymptotic train and test error obtained by fixing $a(t)=\gamma\sqrt{m}$, and not letting second-layer weight evolve. This confirms the hypothesis that $\gamma(t)$ is a slow variabes, while others converge as if $\gamma$ was fixed.
@@ 3 Lower bounding the overfitting timescale
In this section we rigorously establish two results that confirm elements of the scenario outlined in the previous sections. We emphasize that the result presented here are non-asymptotic and do not assume a symmetric initialization of the weights. The first result implies that (under mean field initialization) overfitting cannot take place on times of order one. Our results will be non-asymptotic, i.e. hold at finite $n,m,d$ modulo unspecified absolute constants. For this setting, it is more convenient to rescale time defining $\hat{t}=t\alpha$. In other words, in stead of the flow (1.2), we study
@@
Under the GF dynamics (1.2), and the data distribution in the introduction, further assume $\|\sigma\|_{\mbox{\tiny\rm Lip}},\|\sigma\|_{\infty}\leq L$, $|\varphi(0)|,\|\varphi\|_{\mbox{\tiny\rm Lip}}\leq L$, $\|{\boldsymbol{a}}(0)\|_{\infty}\leq a_{0}$, for some $a_{0}\geq 1$ and that the ${\boldsymbol{w}}_{i}(0)$, $i\leq m$ are independent of the data. Finally assume $n\geq d\vee m$. Then, there exists a universal constant $C_{0}$ and $c=c(L,\tau)>0$ such that, with probability at least $1-2\exp(-cn)$, the following holds for all $\hat{t}\geq 0$,
$\|{\boldsymbol{a}}(\hat{t})\|_{\infty}\leq a_{0}+a_{1}\,\hat{t}\,,\;\;\;\;a_{1}:=C_{0}L(\tau+a_{0}L)\,$, (3.2)
$\mathscrsfs{R}({\boldsymbol{a}}(\hat{t}),\boldsymbol{W}(\hat{t}))-\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}}(\hat{t}),\boldsymbol{W}(\hat{t}))\leq C_{1}L^{3}(a_{0}+a_{1}\hat{t})^{2}\cdot\sqrt{\frac{d}{n}}\,.$ (3.3)
@@
Under the the GF dynamics (1.2), and the data distribution in the introduction, further assume that $\|\varphi\|_{\infty}$, $\|\varphi^{\prime}\|_{\infty}$, $\|\varphi^{\prime}\|_{\mbox{\tiny\rm Lip}}\leq L$, $\|\sigma\|_{\infty}$, $\|\sigma^{\prime}\|_{\infty}$, $\|\sigma^{\prime}\|_{\mbox{\tiny\rm Lip}}\leq L$. Further assume $|a_{i}(0)|\leq L$ for all $i\leq m$, $({\boldsymbol{w}}_{i}(0))_{i\leq m}\sim_{iid}{\sf Unif}({\mathbb{S}}^{d-1})$. Then for any $\delta>0$ there exist constants $c_{0}$ $c_{1}$, $C$ depending on $L,\tau,\delta$ such that, letting $T_{\mbox{\rm\tiny lb}}=c_{0}(\log m)^{1/3}\wedge(\log n/d)^{1/3}$, the following happens with probability at least $1-2\exp(-c_{1}d)$,
@@ 4 Discussion
Interpolation mechanism. In the current setting, the neural model complexity is proportional to $\|{\boldsymbol{a}}(t)\|_{1}/\sqrt{m}=\gamma(t)+o_{n}(1)$. We observe two alternative scenario. If the complexity at initialization is large enough $\gamma_{0}>\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha)\tau$, then the gradient flow rapidly converges to a near interpolator without significant change in $\gamma(t)$. If instead, $\gamma_{0}<\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha)\tau$, then $\gamma(t)$ grows to reach the interpolation threshold at which point the training error converges to 0.
@@
Decoupling of learning and overfitting. When $\gamma_{0}=o_{m}(1)$, the above implies a large-$m$ decoupling between learning (which takes place on faster timescales, as long as $\gamma(t)=o_{m}(1)$), and overfitting (which takes place on slower timescales, when $\gamma(t)=\Omega_{m}(1)$). This has several implications for the questions outlined in the introduction. Question Q3: Lazy initialization $a(0)\asymp\sqrt{m}$ leads to poor generalization because the feature-learning phase is skipped either partially or altogether. Question Q2: Training until interpolation is generally suboptimal. Question Q4: The optimal tradeoff is obtained at the end of the first phase. Further, at fixed overerparametrization $n/md=\alpha$, overfitting starts later for larger models (questions Q5, Q6).
Overfitting and feature unlearning. The above description points at a non-monotonicity of the model quality, which improves on short time scales, and deteriorates at larger time scales. Reciprocally, early stopping acts as a regularization. While this phenomenon is well understood for linear models [FHT00, YRC07], our analysis provides an analogous (quantitative) scenario for training neural network models. In particular, it clarifies the underlying mechanism: in the same dynamical regime in which network complexity grows ($\gamma(t)$ becomes of order one), and training error becomes negligible, the low-dimensional latent features are ‘unlearned’ (${\boldsymbol{v}}(t)$ becomes of order $1/\sqrt{m}$).
@@ Appendix A Setting
${\mathbb{E}}\big{\{}f({\boldsymbol{x}};{\boldsymbol{a}}_{1},\boldsymbol{W}_{1})\,f({\boldsymbol{x}};{\boldsymbol{a}}_{2},\boldsymbol{W}_{2})\big{\}}=\frac{1}{m^{2}}\sum_{i,j=1}^{m}a_{1,i}a_{2,j}h\left(\langle{\boldsymbol{w}}_{1,i},{\boldsymbol{w}}_{2,j}\rangle\right)\,$, (A.4)
where
@@
The Lagrange multipliers $\nu_{i}$ are added to enforce the spherical constraint $\|{\boldsymbol{w}}_{i}(t)\|^{2}=1$. While we consider the case of normalized first-layer weights, our approach can be generalized to unconstrained weights or to include weight decay (ridge regularization). As explained in the main text, we will replace this by gradient flow in the Gaussian model $\widehat{\mathscrsfs{R}}^{g}_{n}({\boldsymbol{a}},\boldsymbol{W})$. We refer to Section J for a discussion of DMFT in the original non-Gaussian model.
@@ Appendix B Technique
$C_{ij}^{n}(t_{1},t_{2})=\langle{\boldsymbol{w}}_{i}(t_{1}),{\boldsymbol{w}}_{j}(t_{2})\rangle\,,\;\;\;\;{\boldsymbol{v}}^{n}_{i}(t):={\boldsymbol{U}}^{{\sf T}}{\boldsymbol{\theta}}_{i}(t)\,,\;\;\;a^{n}_{i}(t)\,.$ (B.4)
@@
Our main focus is on the behavior of the solutions of these equations for large $m$ and, at first sight, the complexity of the DMFT increases with $m$. An important simplification arises when choosing a symmetric initial condition $a_{i}(0)=a_{0}$ for all $i\leq m$, and $({\boldsymbol{w}}_{i}(0))_{i\leq m}\sim_{iid}{\sf Unif}({\mathbb{S}}^{d-1})$. Namely, the solution of the DMFT equations is symmetric under permutations of the neurons: $C_{ii}(t_{1},t_{2})=C_{d}(t_{1},t_{2})$ for $i\leq m$ and $C_{ij}(t_{1},t_{2})=C_{o}(t_{1},t_{2})$ for $i\neq j\leq m$, while ${\boldsymbol{u}}_{i}(t)={\boldsymbol{u}}(t)$, $a_{i}(t)=a(t)$ for $i\leq m$. We then have a reduction to a set of integro-differential equations on $k+3$ functions, that depend parametrically on $m$.
@@
$\lim_{m\to\infty}{\boldsymbol{v}}\Big{(}t_{\#}(m)\cdot\hat{t};m,\overline{\alpha}=\frac{\alpha}{m}\Big{)}={\boldsymbol{v}}_{*}(\hat{t};\alpha)\,.$ (B.5)
@@ Appendix C Dynamical Mean Field Theory (DMFT)
In this section we state the results of Dynamical Mean Field Theory (DMFT). We will outline a heuristic derivation in Section K. We first introduce the general DMFT equations in Section C.1 and the corresponding predictions for certain observable of interest in Section C.2. These are a set of $\Theta(m^{2})$ integro-differential equations in as many unknown functions.

### v3

@@ 2.4 Training on data with latent structure: mean field initialization
Figure 4: Training dynamics under a single-index model. We set $h(q)=\widehat{\varphi}(q)=(9/10)q+q^{3}/6$, $\tau=0.3$ and $\alpha=0.3$, under mean field initialization. Left: second-layer weights. Right: train and test error. Symbols are empirical results for SGD with actual two-layer neural networks with $d=200$, $n=\alpha md$ (averaged over 1010 simulations). Lines correspond to predictions from the DMFT (on the right, continuous: train error; dashed: test error).
We initialize $a(0)=a_{0}$, independent of $m$ and let second layer weights evolve. Note that at initialization the network’s Rademacher complexity is small, namely of order $a_{0}\sqrt{d/n}=a_{0}/\sqrt{\alpha m}$. Our DMFT analyisis predicts two dynamical regimes for large $m$. We will refer to them as ‘first’ and ‘third regime’ for consistency with other settings ( see Sec.G.2 of the appendix). For an illustration, we refer to Figs. 4 and 5.
@@
Namely, $\lim_{m\to\infty}a(t)={a}^{\mbox{\tiny\rm mf1}}(t)$, $\lim_{m\to\infty}{\boldsymbol{v}}(t)={{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}(t)$, and DMFT reduces to a system of $k+1$ ordinary differential equations for the $k+1$ scalar variables $({a}^{\mbox{\tiny\rm mf1}}(t),{{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}(t))$
@@
An explicit formula for ${e}^{\mbox{\tiny\rm mf1}}(t)$ is given in Appendix G.2.1. In the case $k=1$ and $\widehat{\varphi}(z)=h(z)$, we have that ${a}^{\mbox{\tiny\rm mf1}}=1$, ${v}^{\mbox{\tiny\rm mf1}}=1$ is a fixed point of Eq. (2.10), and indeed the only fixed point with ${v}^{\mbox{\tiny\rm mf1}}>0$. If $h^{\prime}(0)>0$, then, we have $({a}^{\mbox{\tiny\rm mf1}}(t),{v}^{\mbox{\tiny\rm mf1}}(t))\to(1,1)$ as $t\to\infty$, and therefore test and train error converge to the Bayes error ${e}^{\mbox{\tiny\rm mf1}}(t)\to\tau^{2}/2$. This is significantly smaller than the test error achieved with lazy initialization. The separation between lazy and mean-field initialization is expected because feature learning takes place in the mean field regime.
Third dynamical regime: $t=\Omega(m)$. Computing the local stability of DMFT solutions around the mean field asymptotics (see Appendix G.2.2) suggests that the latter breaks down for $t=\Theta(m)$. For $t\gtrsim m$, we observe that the second layer weights grow to achieve $a(t)\asymp\sqrt{m}$, the projection onto the latent space decreases to ${\boldsymbol{v}}(t)\asymp 1/\sqrt{m}$, and train and test error diverge, eventually achieving $e_{\mbox{{\rm\footnotesize tr}}}(t)\approx 0$ and test error significantly larger than the Bayes error achieved earlier. We refer to this phenomenon as ‘feature unlearning.’
@@
We observe that the limit curves $(\gamma,{{\varepsilon}}^{\mbox{\tiny\rm mf}}_{\mbox{{\rm\footnotesize tr}}}(\gamma))$, $(\gamma,{{\varepsilon}}^{\mbox{\tiny\rm mf}}_{\mbox{\rm\footnotesize ts}}(\gamma))$, match closely asymptotic train and test error obtained by fixing $a(t)=\gamma\sqrt{m}$, and not letting second-layer weight evolve. This confirms the hypothesis that $\gamma(t)$ is a slow variable, while others converge as if $\gamma$ was fixed.
@@ 3 Lower bounding the overfitting timescale
In this section we rigorously establish two results that confirm elements of the scenario outlined in the previous sections. We emphasize that the result presented here are non-asymptotic, i.e. hold at finite $n,m,d$ modulo unspecified absolute constants. Further, we do not assume a symmetric initialization of the weights. Throughout this section setting, it is more convenient to rescale time defining $\hat{t}=t\alpha$. Hence. instead of the flow (1.2), we study
@@
For $\alpha=\Theta(1)$ the parametrizations $t$ and $\hat{t}$ are equivalent.
The first result of this section implies that (under mean field initialization) overfitting cannot take place on times of order one.
@@
Under the GF dynamics (1.2), and the data distribution in the introduction (with $k$ arbitrary), further assume $\|\sigma\|_{\mbox{\tiny\rm Lip}},\|\sigma\|_{\infty}\leq L$, $|\varphi(0)|,\|\varphi\|_{\mbox{\tiny\rm Lip}}\leq L$, $\|{\boldsymbol{a}}(0)\|_{\infty}\leq a_{0}$, for some $a_{0}\geq 1$ and that the ${\boldsymbol{w}}_{i}(0)$, $i\leq m$ are independent of the data $\{(y_{i},{\boldsymbol{x}}_{i}):i\leq n\}$. Finally assume $n\geq d\vee m$. Then, there exist universal constants $C_{0},C_{1}$, and the following holds for all $\hat{t}\geq 0$,
$\|{\boldsymbol{a}}(\hat{t})\|_{\infty}\leq a_{0}+a_{1}\,\hat{t}\,,\;\;\;\;a_{1}:=C_{0}L(\tau+\sqrt{k}+a_{0}L)\,$, (3.2)
$\big|\mathscrsfs{R}({\boldsymbol{a}}(\hat{t}),\boldsymbol{W}(\hat{t}))-\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}}(\hat{t}),\boldsymbol{W}(\hat{t}))\big|\leq C_{1}(L^{2}(a_{0}+a_{1}\hat{t})^{2}+\tau^{2})\cdot\sqrt{\frac{d}{n}}\,.$ (3.3)
@@
The mean field prediction for test error is the same as for training error and given by
$)\displaystyle e_{\mbox{\rm\footnotesize ts}}(\hat{t})=\frac{1}{2}\|\varphi\|^{2}_{L^{2}}-\frac{1}{m}\sum_{j=1}^{m}{a}^{\mbox{\tiny\rm mf1}}_{j}(\hat{t})\hat{\varphi}({{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}_{j}(\hat{t}))+\frac{1}{2m^{2}}\sum_{j=1}^{m}{a}^{\mbox{\tiny\rm mf1}}_{i}(\hat{t}){a}^{\mbox{\tiny\rm mf1}}_{j}(\hat{t})\,h(\langle{{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}_{i}(\hat{t}),{{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}_{j}(\hat{t}$$\rangle)$
@@
Under the the GF dynamics (1.2), and the data distribution in the introduction (with $k$ arbitrary), further assume that $\|\varphi\|_{\infty}$, $\|\varphi^{\prime}\|_{\infty}$, $\|\varphi^{\prime}\|_{\mbox{\tiny\rm Lip}}\leq L$, $\|\sigma\|_{\infty}$, $\|\sigma^{\prime}\|_{\infty}$, $\|\sigma^{\prime}\|_{\mbox{\tiny\rm Lip}}\leq L$. Further assume $|a_{i}(0)|\leq L$ for all $i\leq m$, $({\boldsymbol{w}}_{i}(0))_{i\leq m}\sim_{iid}{\sf Unif}({\mathbb{S}}^{d-1})$. Then for any $\delta>0$ there exist constants $c_{0}$ $c_{1}$, $C$ depending on $L,\tau,\delta,k$ such that, letting $T_{\mbox{\rm\tiny lb}}=c_{0}(\log m)^{1/3}\wedge(\log n/d)^{1/3}$, the following happens with probability at least $1-2\exp(-c_{1}d)$,
@@ 4 Discussion
Interpolation mechanism. In the current setting, the neural model complexity is proportional to $\|{\boldsymbol{a}}(t)\|_{1}/\sqrt{m}=\gamma(t)+o_{n}(1)$. We observe two alternative scenarios. If the complexity at initialization is large enough $\gamma_{0}>\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha)\tau$, then the gradient flow rapidly converges to a near interpolator without significant change in $\gamma(t)$. If instead, $\gamma_{0}<\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha)\tau$, then $\gamma(t)$ grows to reach the interpolation threshold at which point the training error converges to 0.
@@
Decoupling of learning and overfitting. When $\gamma_{0}=o_{m}(1)$, the fact that $\gamma(t)$ acts as a slow variable implies a large-$m$ decoupling between learning (which takes place on faster timescales, as long as $\gamma(t)=o_{m}(1)$), and overfitting (which takes place on slower timescales, when $\gamma(t)=\Omega_{m}(1)$). This has several implications for the questions outlined in the introduction.
Q3: Lazy initialization $a(0)\asymp\sqrt{m}$ leads to poor generalization because the feature-learning phase is skipped either partially or altogether.
Q2: Training until interpolation is generally suboptimal.
Q4: The optimal tradeoff is obtained at the end of the first phase.
Q5, Q6: Further, at fixed overerparametrization $n/md=\alpha$, overfitting starts later for larger models.
Overfitting and feature unlearning. The above description points at a non-monotonicity of the model quality, which improves on short time scales, and deteriorates at larger time scales. Reciprocally, early stopping acts as a regularization. While this phenomenon is well understood for linear models [FHT00, YRC07], our analysis provides an analogous (quantitative) scenario for training neural network models. In particular, it clarifies the underlying mechanism: in the same dynamical regime in which network complexity grows ($\gamma(t)$ becomes of order one), and training error becomes negligible, the low-dimensional latent features are ‘unlearned’ (${\boldsymbol{v}}(t)$ becomes of order $1/\sqrt{m}$). We expect that these findings also allow to understand the beneficial effect of regularization on the second layer.
@@ Appendix A Setting
${\mathbb{E}}\big\{f({\boldsymbol{x}};{\boldsymbol{a}}_{1},\boldsymbol{W}_{1})\,f({\boldsymbol{x}};{\boldsymbol{a}}_{2},\boldsymbol{W}_{2})\big\}=\frac{1}{m^{2}}\sum_{i,j=1}^{m}a_{1,i}a_{2,j}h\left(\langle{\boldsymbol{w}}_{1,i},{\boldsymbol{w}}_{2,j}\rangle\right)\,.$ (A.4)
The average in the rhs of Eq. (A.4) is over the data distribution while the function $h(q)$ is defined as
@@
The Lagrange multipliers $\nu_{i}$ are added to enforce the spherical constraint $\|{\boldsymbol{w}}_{i}(t)\|^{2}=1$. While we consider the case of normalized first-layer weights, our approach can be generalized to unconstrained weights or to include weight decay (ridge regularization). As explained in the main text, we will replace this by gradient flow in the Gaussian model $\widehat{\mathscrsfs{R}}^{g}_{n}({\boldsymbol{a}},\boldsymbol{W})$. We refer to Section K for a discussion of DMFT in the original non-Gaussian model.
@@ Appendix B Technique
$C_{ij}^{n}(t_{1},t_{2})=\langle{\boldsymbol{w}}_{i}(t_{1}),{\boldsymbol{w}}_{j}(t_{2})\rangle\,,\;\;\;\;{\boldsymbol{v}}^{n}_{i}(t):={\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}_{i}(t)\,,\;\;\;a^{n}_{i}(t)\,.$ (B.4)
@@
Our main focus is on the behavior of the solutions of these equations for large $m$ and, at first sight, the complexity of the DMFT increases with $m$. An important simplification arises when choosing a symmetric initial condition $a_{i}(0)=a_{0}$ for all $i\leq m$, and $({\boldsymbol{w}}_{i}(0))_{i\leq m}\sim_{iid}{\sf Unif}({\mathbb{S}}^{d-1})$. Namely, the solution of the DMFT equations is symmetric under permutations of the neurons: $C_{ii}(t_{1},t_{2})=C_{d}(t_{1},t_{2})$ for $i\leq m$ and $C_{ij}(t_{1},t_{2})=C_{o}(t_{1},t_{2})$ for $i\neq j\leq m$, while ${\boldsymbol{v}}_{i}(t)={\boldsymbol{v}}(t)$, $a_{i}(t)=a(t)$ for $i\leq m$. We then have a reduction to a set of integro-differential equations on $k+3$ functions, that depend parametrically on $m$.
@@
$\lim_{m\to\infty}{\boldsymbol{v}}\Big(t_{\#}(m)\cdot\hat{t};m,\alpha=\frac{\overline{\alpha}}{m}\Big)={\boldsymbol{v}}_{*}(\hat{t};\alpha)\,.$ (B.5)
@@ Appendix C Dynamical Mean Field Theory (DMFT)
In this section we state the results of Dynamical Mean Field Theory (DMFT). We will outline a heuristic derivation in Section L. We first introduce the general DMFT equations in Section C.1 and the corresponding predictions for certain observable of interest in Section C.2. These are a set of $\Theta(m^{2})$ integro-differential equations in as many unknown functions.